# The Ultimate Configuration Management Tool? Lessons from a Mixed Methods Study of Ansible’s Challenges

### v2

@@ 7. Discussion
This section provides a set of actionable insights and discusses the implications of our findings for the IaC ecosystem.
7.1. Actionable Insights
This subsection proposes concrete, actionable recommendations for improving the Ansible ecosystem and the IaC ecosystem in general.
Debugging
As noted in previous studies, debugging in IaC remains a significant challenge (Guerriero et al., 2019). In our analysis of users’ Q&A data, reliability challenges — particularly “Error Handling” — were among the most frequently cited frustrations. Likewise, 10 interview participants described how diagnosing failures becomes particularly difficult when a playbook “gets stuck” with little to no logging feedback. Multiple interviewees lamented the absence of rollback-on-failure or built-in “step-by-step” debugging capabilities that might elucidate the exact point of failure. Although Ansible’s stateless design appeals to many users, it poses challenges for scenarios requiring continuous enforcement or “undo” functionality.
For developers of IaC tools: In DevOps, where observability and rapid troubleshooting are essential, developers should make efforts to provide end users with comprehensive logging. Moreover, clear guidelines, or even mandatory standards, should be established for community-contributed modules to enforce consistent logging practices. IaC tools such as Ansible should develop automatic rollback mechanisms to ensure recovery from failures. Such mechanisms may rely on saving the system’s prior state or executing compensating operations to revert unsuccessful changes. Rollback support might enhance user confidence during IaC script testing and facilitate recovery from unexpected behavior.
For researchers: Participants emphasized the need for better debugging support in IaC tools as IaC tools lack the robust debugging utilities available in traditional languages like C and Python. Although Ansible offers a debugger (Red Hat, 2025a), practitioners reported difficulties and often resorted to shell scripts instead. Additionally, most debugging requires manual intervention, making troubleshooting time-consuming and prone to errors. We propose developing better debugging and automated repair tools that identify common misconfigurations, detect inconsistencies, and automatically suggest or apply fixes. While repair mechanisms exist in traditional software development, their adoption in IaC remains limited (Weiss et al., 2017).
Syntax
Although Ansible’s declarative approach and reliance on YAML/Jinja2 can simplify basic configuration tasks, our analysis of Q&A data found syntax issues with YAML, Jinja and other parts of Ansible were the main challenges of the Usability theme. Additionally, loops and conditional executions accounted for a substantial portion of the reported difficulties. Several interviewees (n = 10) mentioned the limitations in expressing complex logic. They described frustrations with YAML syntax quirks (especially for loops and conditionals), cumbersome variable handling, and a lack of “proper” language features for advanced use cases.
For developers of IaC tools: Beyond the feedback obtained in our study, developers have consistently identified YAML as highly complex and error-prone. This issue is particularly concerning, given that YAML is used not only by Ansible but also by other IaC tools such as Kubernetes. To address this challenge, Asseldonk has proposed replacing YAML with simpler configuration languages, such as TOML (Tom Preston-Werner, 2025). However, to introduce higher-level language concepts, we must go beyond configuration languages. Solutions such as Pulumi (Pulumi, 2025) present an alternative paradigm where IaC is implemented in imperative programming languages rather than configuration languages such as YAML, while still maintaining declarative features. Another example is Nix (NixOS Contributors, 2025), which uses a functional programming language specifically designed for package definitions. Developers of IaC tools must carefully design or select the configuration language, as limitations in syntax and expressiveness can significantly influence the developers’ experience.
Performance
Our topic modeling analysis revealed a substantial volume of posts on slow execution (17.7% of the posts related to Performance and Scalability). In interviews, seven participants reported frustrations with performance bottlenecks, particularly when managing hundreds of hosts. They cited sequential task execution, agentless overhead, and limited parallelization features as key obstacles slowing down their workflows. Participants and our Q&A data analysis converge on the need for better parallelization and reduced overhead in large deployments.
For Ansible developers: Jetporch (Jetporch Contributors, 2025) is a Rust-based Ansible alternative that aims to address performance limitations by avoiding Python (and its inherent execution overhead) and optimizing execution. Jetporch’s design ideas show key areas where Ansible could improve. Ansible should undergo refactoring to introduce parallelization improvements, optimize SSH connections, and enable partial execution in a more efficient manner.
For developers of IaC tools: We recommend treating performance optimization as a priority. Developers should prioritize architectural improvements that enable efficient parallel execution across multiple hosts, optimize communication overhead (particularly for agentless tools relying on SSH), and provide configurable execution strategies for different deployment scales.
Documentation
Documentation problems were mentioned more often in the interviews than in our Q&A data (only 2% of the recent posts). In our survey, eight participants flagged documentation gaps as a pain point. Users commonly struggle to locate authoritative references for advanced features or specialized modules. Interviewees recounted sifting through outdated or misleading community-contributed tutorials.
For Ansible developers: Good documentation is one of the things that our participants liked about Ansible, so we argue that Ansible should capitalize on this fact by improving its current documentation. Beyond traditional methods of documentation overhaul, generative AI tools could dynamically address knowledge gaps. LLMs could identify outdated sections by continuously parsing documentation and flagging inaccurate sections, automatically suggesting updates.
For developers of IaC tools: We recommend prioritizing comprehensive and well-maintained documentation. Developers should establish clear guidelines distinguishing official modules from community contributions, maintain version-specific documentation that is easily navigable, and implement regular audits to identify and update outdated content. Additionally, exploring AI-assisted documentation tools could help maintain accuracy at scale by automatically detecting deprecated information and generating contextual examples for edge cases.
7.2. Future Work and Broad IaC Implications
Our findings highlight key challenges and opportunities for Ansible’s evolution and offer broader insights into the trade-offs inherent in IaC tool design.
Mixed-Methods Framework. We observe differences between publicly discussed topics and the more subjective issues revealed through direct interviews. While thematic areas such as “Control Flow” or “Lack of Support” dominate forum posts, other challenges — like documentation gaps or configuration drift — appear more in interviews. For instance, information gaps and misleading advice may not prompt a direct question on a forum. Practitioners may not even realize they are facing an information gap, nor might they post about it explicitly. Instead, they continue troubleshooting, or they might consult colleagues offline. In interviews, however, participants openly discussed these intangible pain points — highlighting subjective experiences. We argue that qualitative insights should be used with quantitative or text-mining approaches to provide a holistic understanding of IaC adoption barriers. We provide a straightforward and replicable methodology that can be used for other major IaC platforms (e.g., Chef, Puppet) to yield more holistic, data-driven evaluations of challenges that practitioners face.
Balancing Usability and Scalability. User experience seems to vary by expertise level. Beginners emphasize documentation clarity, onboarding simplicity, and accessible debugging tools. Meanwhile, advanced practitioners advocate for performance optimizations (e.g., concurrency) and advanced programming capabilities. This suggests that IaC tools should provide tiered user experiences — offering simplified workflows for beginners while exposing advanced performance optimizations and parallelization capabilities for experts. Striking this balance seems to be important to ensure long-term adoption across skill levels.
Documentation Gaps are a Universal Problem. Finally, documentation gaps and information drift are well-documented challenges across software systems, and our findings confirm their presence in Ansible as well. While well-developed official documentation can lower the barrier to entry (Robillard, 2009; Aghajani et al., 2019), many tools struggle to keep pace with rapid feature expansion and a growing user base (Aghajani et al., 2020). Establishing centralized, consistently updated repositories of best practices and official modules is important for sustained usability.
@@ 8. Threats to Validity
Threats to Internal Validity. A threat arises from the use of LLMs through TopicGPT. LLMs are prone to hallucination and can generate unreliable results, potentially affecting the precision of our findings. To mitigate this, we perform manual validation in two stages: (1) during the generation of the topic list, and (2) during the assignment of topics to individual posts. Also, conducting semi-structured interviews introduces risks such as interviewer bias, where the phrasing or emphasis of questions might influence participant responses. To minimize this, we followed a consistent interview protocol and performed pilots before deploying the study.
Threats to External Validity. Practitioners can discuss Ansible issues on alternative Q&A platforms that we did not analyze, potentially limiting the representativeness of our findings. To mitigate this, we collected posts from three distinct platforms. Another threat is that while some of our findings may generalize to other IaC tools, particularly configuration management tools, more research is needed to confirm their applicability. Future studies should explore other IaC tools to assess the generalizability of our results. Regarding the interviews, we recruited participants primarily through online channels, so our sample may not be representative of the broader Ansible user community. Self-selection bias may have resulted in participants who are more engaged with the tool or have stronger opinions about it. Furthermore, our sample’s distribution across expertise levels may not accurately reflect the actual distribution in the community.
@@ 9. Related Work
To identify challenges that practitioners face in IaC, Rahman et al. analyzed 2,758 Puppet-related questions from Stack Overflow (Rahman et al., 2018). They used qualitative analysis to identify 16 major categories of questions. Many of these categories, including Syntax Error and Filesystem, are also observed in our study. While Rahman et al.’s study focuses on Puppet, our study focuses on Ansible, which, as of 2024, is a more widely used technology for configuration management of services. Furthermore, Rahman et al. analyzed questions posted between January 2010 and December 2016, while our interviews and posts reflect the scenario as of 2024.
Guerriero et al. conducted 44 semi-structured interviews with IaC practitioners to examine adoption and development practices, available tools, and the needs and challenges of working with IaC (Guerriero et al., 2019). Similarly to our study, they highlight issues such as difficulties with testing and debugging. However, we provide a more granular and in-depth analysis, as well as the identification of unique issues of Ansible, such as issues with loops and Jinja.
@@
To the best of our knowledge, the only other mixed methods study related to IaC challenges is the one conducted by Tanzil et al. (Tanzil et al., 2023). However, their study is broader, addressing DevOps challenges in general, with IaC as just one category. They used LDA to perform topic modeling on approximately 174,000 Stack Overflow posts, identifying 23 distinct topics. These were then grouped into four categories, with IaC being one of them, encompassing seven specific subtopics. Finally, the authors validated and extended their findings with a survey of 21 professional DevOps practitioners. Although there is some overlap in the findings between our study and that of Tanzil et al. (e.g., both identify challenges related to File Management and Syntax), we focus exclusively on challenges related to Ansible and IaC, which provides a more focused scope with different insights such as difficulties with loops and conditionals. We also employ a different topic modeling approach and focus on interviews over surveys.

### v3

@@ 7. Discussion
This section discusses the implications of our findings for Ansible as well as for other IaC technologies.
7.1. Failure Locality Guarantees
Our quantitative and qualitative analyses indicate that many Ansible challenges stem from difficulties in localizing failures during playbook execution. High-frequency topics such as Error Handling, Undefined Variables, and YAML, Jinja and Quoting-related syntax issues frequently involve failures that surface far from their root causes. Interview participants consistently described situations where Ansible “gets stuck”, produces misleading errors, or provides insufficient context to determine where execution went wrong. Several interviewees explicitly noted limitations in the current error messages and expressed a need for clearer, more informative feedback. Notably, these observations were made by interviewees with substantial Ansible experience (on average, approximately 4 years), suggesting that these difficulties persist beyond initial onboarding.
We conceptualize this recurring pattern as failure locality: the degree to which an IaC tool enables practitioners to clearly identify the origin and propagation of failures during execution. In Ansible, low failure locality manifests when errors arise late, lack actionable context, or are disconnected from their causes.
This lack of failure locality is reflected in the interviews, where half of the participants reported that Ansible is hard to debug and frequently cited the absence of meaningful progress or termination feedback during execution. As discussed in Section 4, a participant reported a playbook that stalled for 15 minutes despite being expected to run in one minute, with no visible terminal feedback. In such cases, users cannot determine whether the execution is progressing or identify the task responsible for the stall. As a result, failure localization often requires manual interruption, additional logging, or trial-and-error instrumentation.
This lack of execution feedback is also reflected in community discussions. We identified 160 recent posts ($\approx$0.7%) whose titles contain terms such as stuck, hang, or freeze, suggesting that non-terminating or stalled executions are a recurring concern. We restricted the search to post titles to avoid overcounting incidental mentions and to focus on posts where non-termination was the central issue. These reports align with documented Ansible issues that describe stalled executions with limited diagnostic output (Ansible Community, 2016, 2017). Many of these posts were attributed topics related to SSH connection issues (50 posts), AWX (23), and WinRM (10), which is Ansible’s primary remote management protocol for Windows hosts, suggesting that stalled executions often arise at integration boundaries, such as remote connections and AWX-mediated execution orchestration, where Ansible provides limited visibility into progress and blocking conditions.
Insight for Ansible: Ansible could provide phase-aware execution progress and stall diagnostics at integration boundaries such as remote connections. Surfacing the current execution phase (e.g., connection establishment or module execution), along with task and host context, and warning when a phase stalls would help practitioners localize blocking behavior without manual interruption.
The low failure locality in Ansible is not limited to a single class of failures. Our analysis shows that failures related to YAML syntax and quoting exhibit particularly poor locality characteristics, making them among the most prominent usability challenges faced by practitioners. An inspection of the Q&A posts revealed a recurring error message that appears in 706 posts: “The error appears to have been in file X: line Z, column Y, but may be elsewhere in the file depending on the exact syntax problem.”
Such messages acknowledge the uncertainty about the fault’s location, forcing users to search beyond the reported location. This behavior undermines failure locality, as the system is unable to confine the failure to the construct that caused it. Instead of narrowing the search space, the diagnostic output expands it, increasing the cognitive effort required to diagnose even minor errors.
The interview data further illustrate the practical impact of this loss of locality. As mentioned in Section 4, a participant with 9 years of Ansible experience spent several hours debugging a YAML quoting issue. This example highlights how small, localized mistakes can propagate into failures that are difficult to attribute, diagnose, and correct due to ambiguous or misleading error reporting.
Insight for YAML-based IaC technologies: Improving debugging effectiveness requires compensating for YAML’s weak syntactic failure locality. YAML-based IaC tools could provide early, schema-guided validation, precise diagnostics for quoting and escaping errors that identify the exact responsible construct instead of approximate file locations, and clear distinctions between parsing, templating, and execution failures, thereby reducing the diagnostic search space.
7.2. Blurred Language Boundaries
Our Q&A analysis indicates that issues related to Loops, Jinja2, Conditional Execution, and Undefined Variables are among the most frequently reported by practitioners Moreover, these topics exhibit strong co-occurrence in Ansible-related discussions. Table 1 reports the pairwise co-occurrences and coverage among these four topics. Additionally, across all topics, roughly one-fifth of posts associated with any topic also involve at least one of the others. These results suggest that practitioners reason about these concepts together. They experience them as a tightly coupled cluster of challenges that emerge at the intersection of Ansible’s execution model and the Jinja2 templating system. This is consistent with our interview data, in which participants repeatedly highlighted difficulties with Jinja2, mainly experienced practitioners with an average of seven years of Ansible experience, and suggested the addition of higher-level language constructs as a potential improvement.
Table 1. Pairwise co-occurrences between topics Jinja2, Loops, Conditional Execution, and Undefined Variables.
Coverage (%)
Topic Pair (T1) Topic 2 (T2) Count T1 T2
Jinja2 Loops 134 10.1 9.0
Cond. Exec. Loops 129 11.4 8.7
Jinja2 Undef. Vars. 95 7.2 8.5
Cond. Exec. Undef. Vars. 79 7.0 7.1
Loops Undef. Vars. 76 5.1 6.8
Cond. Exec. Jinja2 67 5.9 5.1
Playbooks operate at the intersection of multiple languages and abstraction layers: YAML as a data serialization format, Ansible-specific schema extensions (e.g. when), and Jinja2 as an embedded templating language. Their composition creates blurred semantic boundaries that are difficult for practitioners to internalize.
This interaction helps to explain an apparent tension in our findings. In RQ2, participants frequently described Ansible’s readable syntax as a key factor influencing adoption. However, our results indicate that this readability may not scale uniformly with the complexity of playbooks. As playbooks become increasingly entangled across language boundaries, practitioners must simultaneously determine in which language they are currently writing in, which evaluation model applies, and where expressions are interpreted. Maintaining an accurate mental model of these shifting boundaries imposes considerable cognitive overhead.
Nested Loops
A clear manifestation of this boundary breakdown appears in the handling of nested loops, which account for 7.3% of all loop-related discussions. Notably, Ansible’s official documentation explicitly discourages nested loops (Red Hat, 2026b): “The simplest way to ‘nest’ loops is to avoid nesting loops, just format the data to achieve the same result”. Instead, Ansible recommends that users restructure data and rely on Jinja2 filters, to simulate nested iteration. This design choice introduces a tension between Ansible’s declarative model and the imperative nature of iteration. While loops are exposed as imperative constructs, their expressiveness is deliberately constrained, pushing users toward a functional programming style implemented via Jinja2 filters. As a result, users are expected to translate imperative intent into functional transformations, increasing cognitive load and reducing readability.
Ansible-specific filters
This tension is further amplified by Ansible’s reliance on Ansible-specific Jinja2 extensions and filters. For example, the documentation prescribes the use of the dict2items filter, an Ansible-specific addition, to iterate over dictionaries (Red Hat, 2026a). We obtained the list of Ansible-specific filters from the official documentation (Ansible, Inc., 2026) and subsequently searched the Q&A posts for corresponding invocation patterns for each filter. We identified the usage of Ansible-specific filters in user-provided examples in 3,014 posts, corresponding to 12.5% of all recent posts.
We then assessed whether these filters were central to the issues discussed in the posts. To this end, we selected posts that not only matched the aforementioned pattern but also referenced the same filter, or a semantically related term, in the post title. For example, for the regex_findall filter, we additionally searched for the terms regex, regular, expression, and findall. Of the 3,014 posts containing filter usages, 706 satisfied this criterion, indicating that at least 3% of all recent posts are related to the use of Ansible-specific filters.
Although technically effective, such constructs deepen the entanglement between Ansible and Jinja2, forcing users to learn not only the Jinja2 semantics but also the customized Ansible dialect. Rather than cleanly separating concerns, this approach layers declarative configuration, imperative execution, and functional transformation into a single syntactic space, making it harder for users to predict behavior and diagnose errors.
Templating boundaries
Inconsistencies in templating boundaries further obscure the mental model practitioners must maintain. Conditional execution (when) implicitly evaluates Jinja2 expressions without requiring the explicit Jinja2 delimiters used in Ansible, while loops require Jinja2 expressions to be explicitly enclosed within these delimiters. In fact, on April 14, 2025, a commit (Ansible Project, 2026) was pushed to Ansible that introduces a new deprecation warning that “Conditionals should not be surrounded by templating delimiters such as {{ }} or {% %}. This feature will be removed from ansible-core version 2.23”. A similar warning had already been introduced in Ansible 2.3, as one collected post mentions (PhilEv and Ansible Community, 2017). The recurrence of this guidance across versions suggests a long-standing source of confusion. Consistent with this, we observe that this warning message appears 37 times in recent posts. Such inconsistencies make it unclear when users are writing Jinja2, how expressions are parsed and evaluated, and consequently weaken boundary visibility.
Taken together, these findings suggest that many of Ansible’s usability challenges may not be due to individual features, but to boundary friction between languages and paradigms that overlap.
Insight for Ansible: Our findings suggest the need to revisit the design and exposure of control-flow abstractions, with a focus on making language boundaries explicit and reducing cross-paradigm leakage. Ansible could benefit from first-class expressive control-flow constructs that align with the mental models of practitioners, including native support for nested iteration and richer conditionals, thus reducing the dependency on Jinja2 as a workaround language and lowering the cognitive overhead introduced by functional-style filters. In parallel, Ansible could strive for consistent templating semantics by standardizing expression evaluation rules across constructs or explicitly signaling evaluation contexts. Treating Jinja2 as a clearly delineated component, with explicit role, scope, and limitations in both syntax and documentation, would improve learnability and reduce boundary confusion.
Insight for designers of new IaC technologies: Our results highlight the risks of polyglot configuration models that blend declarative, imperative, and templating paradigms without clear separation. Our findings show that boundary ambiguity can become a primary source of user difficulty. Future IaC tools could therefore treat boundary design as a first-class concern, explicitly modeling where data declaration ends, where control flow begins, and how (or whether) templating is allowed to intervene.
7.3. Documentation Gaps
Documentation is the primary means by which users learn and troubleshoot IaC tools, however, our findings show that Ansible’s documentation exhibits both explicit gaps (directly reported as missing or inadequate) and implicit gaps, inferred from recurring questions that existing documentation fails to resolve.
Topic modeling indicates that documentation issues constitute a subset of the Information Gaps theme (480 posts,  2% of recent posts). Participants across experience levels (n=10) repeatedly called to improve documentation. E4 described the current official documentation as scattered, noting that “strong concepts are scattered over different pages.” E8 found the documentation examples helpful but noted that “sometimes they don’t [provide] all the examples”
Because topic modeling captures only explicit discussion of documentation, we conducted a post-hoc analysis to identify documentation issues embedded in technical posts. This analysis shows that 5.4% of all posts (1,301) report documentation problems (2.5× higher than topic modeling suggests). This result comes from a two-stage process: a keyword search followed by manual validation. The search identified 1,858 posts (7.7%) mentioning “documentation” or “docs”, which were then refined through review to isolate genuine documentation issues. Using a conservative proportion of 0.5 with finite-population correction, we sampled 353 posts to achieve a 95% CI with a $\pm 5\%$ error margin. One author coded the posts, with a second author independently reviewing the classifications as either true positives (e.g., missing, unclear, or mismatched information) or false positives. Of the sample, 247 posts (70.0% $\pm$ 5%, 95% CI) were true positives and 106 (30.0% $\pm$ 5%, 95% CI) false positives, thus, about $5.05\%$ of recent posts genuinely report documentation issues. We also saw that the percentage of posts reporting documentation problems has increased over time from approximately 4.2% in 2016-2020 to 7.6% in 2024 and 10.9% in early 2025. Because this post-hoc analysis relies on manual interpretation of discussion content, the resulting proportion should be interpreted as an approximate estimate rather than an exact prevalence. Documentation concerns also arise while discussing other technical topics, including Module Usage ($\sim$80 posts), Integration Issues ($\sim$74), Python Issues ($\sim$97), and Error Handling ($\sim$53). Users in posts report missing documentation (“Couldn’t find Ansible documentation on GPG” (Stack Overflow Community, 2019)), mismatches between documentation and behaviour (“If I read the documentation to the letter it should fail” (Ansible Community, 2024)), and finding existing documentation insufficient (“the documentation is still very vague about how this actually works and behaves” (Reddit Ansible Community, 2023)).
Insight for IaC tools: The finding that 5.4% of recent posts report documentation issues within technical discussions suggests that documentation gaps surface during implementation rather than through explicit documentation queries. Although high-quality documentation lowers barriers to entry (Robillard, 2009; Aghajani et al., 2019), many tools fail to keep pace with rapid feature growth and expanding user bases (Aghajani et al., 2020). Tool developers can apply our topic modeling approach to mine community Q&A data and prioritize documentation improvements by technical domain, focusing on topics that most frequently co-occur with documentation mentions. While demonstrated on Ansible, this method generalizes to other IaC and software tools.
7.4. Execution Backends and Scalability
Performance and scalability concerns are reflected in both our quantitative and qualitative results. Performance and Scalability account for 8.1% of recent posts, and within this category, posts explicitly classified under Performance represent 16.3%. An analysis of topic co-occurrence within Performance posts shows that AWX (10.6%), Loops (10.2%), SSH Connection Issues (9.0%), Error Handling (5.6%), and Asynchronous Execution (5.0%) are the most frequent co-occurring topics, where percentages denote the proportion of Performance posts in which each topic appears. The prominence of SSH Connection Issues and Asynchronous Execution indicates that performance problems are frequently associated with remote execution overhead and limitations in parallel execution models.
Interview participants emphasized slow execution and the need to improve speed and parallelization, particularly when managing large infrastructures. Several participants identified the agentless execution model as a contributing factor to scalability limitations, with one E2 describing slow runs as “an eternal pain point”.
This concern must be viewed in light of the factors driving Ansible’s adoption. Participants highlighted the agentless, SSH-based execution model as a key advantage. Rather than suggesting that agentless execution should be abandoned, our findings point to the need for execution backends that preserve this model while mitigating its performance costs. In this context, an experienced participant advocated for integrating solutions similar to Mitogen (Mitogen Contributors, 2025) to reduce SSH overhead and improve execution efficiency.
Mitogen is an alternative execution framework for Ansible that optimizes remote task execution by minimizing repeated SSH connection setup and Python interpreter initialization. Instead of spawning a new remote process for each task invocation, Mitogen establishes persistent execution contexts and reuses them across tasks, significantly reducing communication and startup overhead, while preserving Ansible’s agentless execution model.
Insight for Ansible: Mitogen illustrates that Ansible’s performance limitations are not inherent to declarative IaC workflows, but are shaped by the characteristics of its execution backend. An implication for Ansible is therefore to treat optimized execution frameworks, such as Mitogen-like approaches, as first-class supported components with clear integration into the platform. Such support would allow users to systematically address slow execution and parallelization challenges in large-scale deployments, rather than relying on external or ad hoc optimizations.
7.5. Future Work
We identify opportunities for Ansible’s evolution and broader trade-offs in IaC tool design. First, our concept of failure locality warrants engineering validation through the implementation and evaluation of phase-aware diagnostics and error reporting in controlled experiments. The blurred language boundaries we identified also suggest a need for design guidelines, including formal design patterns for composing declarative, imperative, and templating paradigms in IaC languages, potentially via prototype DSL development. Mining large-scale Ansible repositories to quantify the prevalence of identified anti-patterns (nested loops, Jinja2 complexity, variable scoping issues) would provide complementary behavioral evidence and enable correlation with code quality metrics. We also argue that our mixed methodology of combining large-scale topic modeling of community discussions with targeted practitioner interviews to identify challenges is easily replicable in other contexts. Applying this methodology to tools such as Terraform or Kubernetes could enable a cross-tool comparison of challenge patterns and inform tool selection and improvement priorities across the IaC ecosystem.
@@ 8. Threats to Validity
Threats to Internal Validity.
The use of LLMs via TopicGPT introduces a risk of hallucination and unreliable outputs, potentially affecting result precision. We mitigated this through a two-stage manual validation during topic generation and topic assignment. Semi-structured interviews also risk interviewer bias; this was mitigated by using a consistent protocol and piloting the interviews.
Threats to External Validity.
Our findings may be limited by practitioners discussing Ansible on Q&A platforms not included in our dataset; we mitigated this by collecting data from three distinct platforms. While some findings may generalize to other IaC tools, particularly configuration management tools, further research is needed to confirm their applicability. Interview participants were primarily recruited online, introducing potential self-selection bias and limiting representativeness across expertise levels. As with qualitative research generally, interview results provide contextual insights but are not statistically representative of the broader Ansible community. Our analysis relies on public Q&A platforms whose usage patterns are changing as generative AI adoption increases (Burtch et al., 2024). We mitigated this by restricting analysis to posts from the five years prior to data collection; the resulting corpus of 24,132 posts still exhibits diverse topics, indicating continued platform relevance. Nevertheless, as generative AI increasingly addresses common issues, future forum discussions may shift toward more complex problems, potentially affecting long-term generalizability.
@@ 9. Related Work
To identify challenges that practitioners face in IaC, Rahman et al. analyzed 2,758 Puppet-related questions from Stack Overflow (Rahman et al., 2018), identifying 16 major categories, many of which (e.g., Syntax Error, Filesystem) also appear in our study. While their focus is Puppet, we study Ansible, a more widely used configuration management tool as of 2024, and consider more recent posts and interviews. Furthermore, Rahman et al. analyzed questions posted between January 2010 and December 2016, while our interviews and posts reflect the scenario as of 2024.
Guerriero et al. conducted 44 semi-structured interviews with practitioners to examine IaC practices, tools, and challenges (Guerriero et al., 2019), highlighting issues such as testing and debugging. Our study provides a more granular analysis and identifies Ansible-specific issues such as lack of failure locality and blurred language boundaries.
@@
To our knowledge, the only other mixed-methods study on IaC challenges is that by Tanzil et al. (Tanzil et al., 2023). However, their study addresses DevOps challenges in general, with IaC as just one category. They used LDA to perform topic modeling on approximately 174,000 Stack Overflow posts, identifying 23 distinct topics. These were grouped into four categories, with IaC being one of them, encompassing seven specific subtopics. Finally, the authors validated and extended their findings with a survey of 21 professional DevOps practitioners. Although there is some overlap in the findings between our study and that of Tanzil et al. (e.g., both identify challenges related to File Management and Syntax), we focus exclusively on challenges related to Ansible and IaC, which provides a more focused scope with different insights such as difficulties with YAML and the risks of designing a polygot technology. We also employ a different topic modeling approach and focus on interviews over surveys.